write the statement using absolute value notation.
1) The distance between y and 2 is greater than or equal to 4.

2) The distance between x and c is less than d units
3)y is more than d units from c.

Answer by stanbon(75887)

write the statement using absolute value notation.
1) The distance between y and 2 is greater than or equal to 4.
|y-2| >= 4
y-2 <= -4 or y-2 >=4
y <= -2 or y >= 6
--------------------
2) The distance between x and c is less than d units
|x - c| < d
-----
-d < x-c < d
-----
c-d < x < c+d
------------------------------
3)y is more than d units from c.
|c - y| > d
---
c-y < -d or c-y > d
y > c+d or y < c-d
----------------------
Cheers,
Stan H.
------------------------------